OpenStudy (anonymous):

Who is good at 8th grade math? I will give metals!!!

OpenStudy (anonymous):

Me

OpenStudy (matlee):

yoyoyoyoyoyoyo right here

OpenStudy (matlee):

the answer is A

OpenStudy (anonymous):

A set of equations is given below: Equation M: y = 3x + 4 Equation P: y = 3x + 7 Which of the following options is true about the solution to the given set of equations? (4 points) Two solutions Infinite solutions No solution One solution

OpenStudy (matlee):

ists B

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Solve the system of equations and choose the correct answer from the list of options. d + e = 12 –d + e = –8 Label the ordered pair as (d, e). (4 points) (10, 2) (12, 0) (0, 0) (2, 10)

OpenStudy (amorfide):

one of the equations are in the form y=mx+c substitute the given, into the second equation multiply out of the brackets and simplify to work out x then you can substitute your x value into an original equation and work out y

OpenStudy (anonymous):

@Daniel13519 and @matlee it is against the code of conduct to give out the answer out of bat, you are supposed to help by solving the problem BEFORE you give out the answers ;)

OpenStudy (anonymous):

I have 5 questions

OpenStudy (anonymous):

My bad bro

OpenStudy (yanasidlinskiy):

Direct Answers are against the Code-of-Conduct. Please be sure to review the Conduct before answering someones question. Thank you. Sorry to interrupt you guys. I can report you, but I will not do it because I'm that nice=) Thank you! Ambassador of OpenStudy YanaSidlinskiy

OpenStudy (anonymous):

Already beat you to it @YanaSidlinskiy XD

OpenStudy (anonymous):

I dont think its a big deal tho but i wont do it any more

OpenStudy (anonymous):

@Daniel13519 it is a big deal because, when you give people the answer, that you may guessed wrong and they can get bad grades because of someone guessing answers to them. You may not guess the answer, but it still required to give example and though of how you get that answers. ;)

OpenStudy (yanasidlinskiy):

XD @mitchal lol. Are we really racing? It is a big deal because others will think you are cheating. You can get a warning or a suspension automatically because moderators suspend those whos smartscores are above 50+ and they think they know the rules pretty well. So that's why I don't recommend you doing it and thank you again for not doing it again:)! Sorry to interrupt again. I appreciate it.

OpenStudy (anonymous):

oh that means i need to keep my eye out since im 56 O_O Mitchal soon to be ambassador

OpenStudy (yanasidlinskiy):

Yes you do. In fact, we all need to.